How old is a skeleton that has lost 75% of its Carbon-14? I have NO idea how to do this problem. Could someone help me?
Answer by stanbon(75887)

The half-life of Carbon-14 is 5700 years
The formula you need is
Amount you now have= (Amount you started with)(1/2)^(t/5700)
You want the "Amount you now have" to be 75% of "Amount you started with".
EQUATIOn:
Let the original amount be "x".
0.75x=x(1/2)^(t/5700) where t is the number of years
0.75=(1/2)^(t/5700)
Take the natural log of both sides to get:
ln(0.75)=(t/500)[ln(1/2)]
t/5700= [ln0.75]/[ln(1/2)]
t/5700=0.415037...
t=2365.71 years.
Cheers,
Stan H.